### v2

# EchoTrail-GUI: Building Actionable Memory for GUI Agents via Critic-Guided Self-Exploration

Abstract
Contemporary GUI agents, while increasingly capable due to advances in Large Vision-Language Models (VLMs), often operate with a critical limitation: they treat each task in isolation, lacking a mechanism to systematically learn from past successes. This “digital amnesia” results in sub-optimal performance, repeated errors, and poor generalization to novel challenges. To bridge this gap, we introduce EchoTrail-GUI, a novel framework designed to mimic human-like experiential learning by equipping agents with a dynamic, accessible memory. Our framework operates in three distinct stages. First, during Experience Exploration, an agent autonomously interacts with GUI environments to build a curated database of successful task trajectories, validated by a reward model. Crucially, the entire knowledge base construction is thus fully automated, requiring no human supervision. Second, in the Memory Injection stage, upon receiving a new task, our system efficiently retrieves the most relevant past trajectories to serve as actionable “memories.” Finally, during GUI Task Inference, these memories are injected as in-context guidance to inform the agent’s reasoning and decision-making process. We demonstrate the efficacy of our approach on benchmarks including Android World and AndroidLab. The results show that EchoTrail-GUI significantly improves the task success rate and operational efficiency of baseline agents, validating the power of structured memory in creating more robust and intelligent GUI automation.
1 Introduction
The rapid advancement of Large Vision-Language Models (VLMs) [3, 23] has fundamentally reshaped the development of autonomous GUI agents. Early approaches relied on structured metadata such as accessibility trees [26], whereas modern systems can now parse raw GUI screenshots and follow complex user instructions through sequences of fine-grained actions, including clicking, scrolling, typing, and navigation [7, 15, 32]. These capabilities have enabled significant progress in digital task automation, intelligent computer-use assistants, and interactive multimodal copilots.
Despite these advances, existing GUI agents still suffer from a core limitation that we term digital amnesia. Most current agents remain stateless: they approach each task independently, lacking any mechanism to accumulate operational knowledge or capitalize on prior successful behaviors. This deficiency restricts their robustness, leads to repeated mistakes, and undermines generalization when encountering diverse application layouts or multi-step tasks. Two fundamental obstacles contribute to this limitation.
The first is the experience acquisition bottleneck. High-quality trajectory data is crucial for building generalizable GUI agents [16, 30], yet obtaining such data is expensive. Manual annotation is time-consuming and unscalable, while unguided exploration often produces incoherent or low-utility interaction traces. Recent efforts attempt to synthesize trajectories from videos [9], web tutorials [31, 36], or reverse task inference [18]. Others explore autonomous exploration [28], but typically lack effective mechanisms to control trajectory quality or prevent the accumulation of noisy behaviors.
The second challenge is the knowledge application gap. Even when trajectory corpora are available, efficiently leveraging them remains nontrivial. Although VLMs possess strong in-context learning abilities, most GUI agents rely on static examples or handcrafted prompts [33, 34], failing to retrieve and apply past experience dynamically. Retrieval-augmented methods [29] offer a promising direction, yet their effectiveness hinges on the availability of well-curated, semantically aligned operational memories—resources that are typically scarce or costly to obtain.
To address these challenges, we introduce EchoTrail-GUI, a unified framework that emulates the human-like cognitive cycle of learn–remember–apply. Our central goal is to construct a self-improving, closed-loop system where the agent autonomously gathers experience, distills actionable knowledge, and reuses it to guide new tasks. Inspired by retrieval-augmented generation (RAG) [11], EchoTrail-GUI enables GUI agents to dynamically access the most relevant and successful past trajectories as structured, contextual guidance.
As illustrated in Figure 1, EchoTrail-GUI consists of three stages:
  1. 1.
    Experience Exploration. An exploration agent autonomously interacts with the GUI environment to generate diverse task trajectories. A critic-style reward model evaluates trajectory quality, allowing the system to construct a high-fidelity, human-free repository of successful trajectories.
  2. 2.
    Memory Injection. Given a new task instruction, we retrieve semantically relevant past experiences using a hybrid dense–sparse retrieval strategy. These trajectories serve as contextual memory that reflects the agent’s own operational history.
  3. 3.
    GUI Task Inference. Retrieved trajectories are reformatted into structured guidance and injected into the agent’s prompt, acting as an operational blueprint that supports reasoning, reduces redundant exploration, and improves action prediction.
By integrating these stages into a fully automated pipeline, EchoTrail-GUI transforms stateless GUI agents into memory-augmented systems that grow more capable over time.
In summary, our contributions are as follows:
  • We present EchoTrail-GUI, a three-stage framework that autonomously constructs and leverages actionable memory for GUI agents through critic-guided self-exploration, dynamic retrieval, and memory-augmented inference.
  • We build EchoTrail-4K, a curated dataset containing over 4,000 high-quality trajectories collected in Android environments, supporting both retrieval-based inference and downstream training.
  • Through extensive experiments on AndroidWorld [16] and AndroidLab [30], we demonstrate that memory-augmented GUI agents significantly outperform their stateless counterparts, yielding substantial gains in success rate, sub-goal completion, and operational robustness.
2 Related Work
GUI agents are designed to perceive and interact with digital environments by processing visual screen information. The emergence of powerful Multimodal Large Language Models (MLLMs) [1, 3, 2] has fundamentally shifted the paradigm from relying on structured metadata (like accessibility trees or XML) to end-to-end approaches that directly process raw screen pixels [16, 25, 10]. Research in this area is primarily explored along two vectors. One major research thrust focuses on enhancing the intrinsic capabilities of a single powerful agent. This involves extensive pre-training on large, GUI-centric datasets [15] or proposing unified action spaces to improve interaction abilities [27]. A parallel, training-free strategy leverages the in-context learning and reasoning of MLLMs, often combining reasoning approaches like ReAct [35] with visual annotation techniques [16, 33]. The second vector explores the efficacy of collaborative multi-agent systems. These frameworks typically employ an orchestrator-agent architecture to decompose complex tasks and delegate sub-tasks [5, 10]. In the GUI domain, this has been extended to integrating specialized modules such as reflection agents for performance evaluation and error correction [21], or self-evolving modules that utilize web knowledge and episodic memory for adaptation [37].
The creation of high-fidelity trajectory data is a critical prerequisite for advancing GUI agents, as this data underpins their ability to plan and execute tasks in complex digital worlds [14]. The challenges associated with manual annotation have accelerated research into automated data generation. One approach involves synthesizing trajectories from existing human-generated content. This includes extracting (frame, action) sequences from online videos [9] or converting web tutorials into structured, synthetic trajectories suitable for agent training [31, 36]. A distinct method empowers agents to generate trajectories by autonomously interacting with the GUI environment. This ranges from simple random or seeded exploration [37, 26] to more sophisticated techniques like ”reverse task synthesis,” which infers meaningful high-level tasks from observed low-level interaction traces [18].
Moving beyond using external content solely for training, another paradigm treats these resources as dynamic, non-parametric sources of knowledge accessible at runtime. This Retrieval-Augmented Generation (RAG) approach aims to enhance agent performance without necessitating model fine-tuning. For example, some work employs lightweight models to retrieve relevant guides from a knowledge base, which are then summarized to instruct a frozen agent [29]. Other research focuses on mining ”transition-aware knowledge” from autonomous exploration traces and formatting this knowledge into goal-oriented prompts to guide the agent’s decision-making during execution [28].
3 EchoTrail-GUI
To address the challenges of experience acquisition and knowledge application in GUI agents, we introduce EchoTrail-GUI. Our framework establishes a virtuous framework of autonomous exploration, structured memory consolidation, and experience-guided inference. It operates in three core stages, as illustrated in Figure 1: (I) Critic-Guided Self-Exploration to build a memory base, (II) Dynamic Memory Injection for task-relevant retrieval, and (III) Memory-Augmented Inference for execution.
Figure 1: The Architecture of EchoTrail-GUI. Our framework consists of three stages. (I) Critic-Guided Self-Exploration: An exploration agent $\pi_{\text{explore}}$ generates trajectories ($\tau$) which are evaluated by a critic ($R_{\text{critic}}$). In-progress trajectories guide exploration via a processing database ($D_{\text{proc}}$), while high-quality completed trajectories are archived into a permanent memory database ($D_{\text{mem}}$). (II) Dynamic Memory Injection: For a new task instruction $I$, a hybrid retrieval system fetches the top-$K$ relevant trajectories $M_{t}$ from $D_{\text{mem}}$ to serve as contextual memory. (III) Memory-Augmented Inference: The retrieved memories $M_{t}$ are injected into the prompt of a GUI agent $\pi_{\text{aug}}$, guiding its iterative process of thought, action, and observation to solve the task.
3.1 Problem Formulation
We formulate GUI task automation as a Partially Observable Markov Decision Process (POMDP). At each step $t$, the agent observes the current state $s_{t}$ (a GUI screenshot) and must select an action $a_{t}$ from a discrete action space $\mathcal{A}$ (e.g., click, type). Given a task instruction $I$, the objective is to learn a policy $\pi(a_{t}|s_{t},I)$ that generates a trajectory of state-action pairs $\tau=(s_{0},a_{0},s_{1},a_{1},\dots)$ to fulfill $I$.
Standard agents rely on a frozen, pre-trained policy $\pi_{\text{base}}(a_{t}|s_{t},I,H_{t})$, where $H_{t}$ is the action history. We propose to augment this with a memory component. Our framework first constructs a memory database of successful trajectories, $D_{\text{mem}}$. At inference time, given a new instruction $I$, we retrieve a relevant memory set $M_{t}\subset D_{\text{mem}}$. The agent’s action is then sampled from a memory-augmented policy $\pi_{\text{aug}}$:
$a_{t}\sim\pi_{\text{aug}}(a_{t}|s_{t},I,H_{t},M_{t}).$ (1)
The central challenge, which our framework addresses, is the automated construction of a high-quality $D_{\text{mem}}$ and the effective utilization of $M_{t}$ to guide $\pi_{\text{aug}}$.
3.2 Stage I: Critic-Guided Self-Exploration
The first stage autonomously populates the memory database $D_{\text{mem}}$, a process detailed in Algorithm 1. This stage leverages an exploration agent, a critic model, and a dual-database system to generate and filter high-quality experiences without human intervention.
3.2.1 Autonomous Trajectory Generation
An exploration agent, $\pi_{\text{explore}}$, built upon a foundation VLM, is responsible for generating trajectories. At each step $t$, it takes as input the current screenshot $s_{t}$, its action history $H_{t}$, and real-time guidance $G_{t}$ retrieved from a temporary database (see Sec 3.2.3). It then outputs an action $a_{t}=\pi_{\text{explore}}(s_{t},H_{t},G_{t})$.
To balance breadth of discovery with depth of purpose, we introduce a Progressive Intent Focus mechanism. The agent initiates exploration in a Curiosity-Driven Mode, prioritizing interaction with novel UI elements. After a few steps ($t>T_{\text{focus}}$), it transitions to a Target-Focused Mode, where it first formulates a concrete sub-goal (e.g., ”add item to cart”) and then executes actions to achieve it. This ensures that generated trajectories are both diverse and coherent.
3.2.2 Trajectory Quality Assessment via Critic
Upon termination, each generated trajectory $\tau$ is evaluated by a critic, implemented as a reward model $R_{\text{critic}}$. The critic acts as a quality filter, assigning a score based on the trajectory’s coherence, efficiency, and success in achieving its implicit goal. This score is mapped to a 5-level scale. We define a quality threshold $\theta_{\text{good}}=4$. A trajectory is deemed high-quality and worthy of inclusion in the permanent memory if its score satisfies this condition:
$R_{\text{critic}}(\tau)\geq\theta_{\text{good}}.$ (2)
Importantly, the stored trajectories are not raw screenshot sequences. Instead, to enhance generality, reduce redundancy, and maintain a lightweight memory footprint, each trajectory is converted into an abstracted representation that contains (i) a textual description of the screenshot at each step, (ii) the exploration agent’s intent summary, and (iii) the executed action. This structured, modality-agnostic format preserves the essential semantic and operational information needed for downstream retrieval and in-context guidance, while avoiding the storage cost and device-specific bias inherent in pixel-level data.
This critical filtering step is paramount, as our ablations (Sec 4.3.1) show that injecting low-quality memories is more detrimental than providing no memory at all.
Algorithm 1 Critic-Guided Self-Exploration
1:Input: Exploration agent $\pi_{\text{explore}}$, Critic $R_{\text{critic}}$, Threshold $\theta_{\text{good}}$
2:Initialize: Processing DB $D_{\text{proc}}\leftarrow\emptyset$, Memory DB $D_{\text{mem}}\leftarrow\emptyset$
3:for episode $i=1,\dots,N$ do
4:  $\tau_{\text{new}}\leftarrow(s_{0})$ where $s_{0}$ is the initial state.
5:  for step $t=0,\dots,T_{\text{max}}-1$ do
6:   $H_{t}\leftarrow$ actions in $\tau_{\text{new}}$
7:   $G_{t}\leftarrow\text{RetrieveGuidance}(D_{\text{proc}},\tau_{\text{new}})$
8:   $a_{t}\leftarrow\pi_{\text{explore}}(s_{t},H_{t},G_{t})$
9:   Execute $a_{t}$, observe next state $s_{t+1}$
10:   $\tau_{\text{new}}\leftarrow\tau_{\text{new}}\circ(a_{t},s_{t+1})$ $\triangleright$ Append to trajectory
11:   Update $D_{\text{proc}}$ with the partial trajectory $\tau_{\text{new}}$
12:   if stopping condition is met then break
13:   end if
14:  end for
15:  if $R_{\text{critic}}(\tau_{\text{new}})\geq\theta_{\text{good}}$ then
16:   $D_{\text{mem}}\leftarrow D_{\text{mem}}\cup\{\tau_{\text{new}}\}$ $\triangleright$ Archive high-quality memory
17:  end if
18:end for
19:Return: Memory Database $D_{\text{mem}}$
3.2.3 Dual-Memory Learning System
The exploration process is scaffolded by a dual-database system to enable efficient learning:
  • Processing Database ($D_{\text{proc}}$): A short-term, volatile memory storing both successful and failed in-progress trajectories. Its purpose is to provide immediate, real-time feedback ($G_{t}$) to $\pi_{\text{explore}}$, helping it avoid repeating recent mistakes and reinforcing promising short-term strategies.
  • Memory Database ($D_{\text{mem}}$): The final, curated, and permanent knowledge base. It contains only complete, high-quality trajectories that have passed the critic’s filter. This database is the primary asset for guiding task execution in the subsequent stages.
3.3 Stage II: Dynamic Memory Injection
This stage bridges the gap between past experience and current tasks. Given a new user instruction $I$, we retrieve the $K$ most relevant trajectories, $M_{t}=\{\tau_{1},\dots,\tau_{K}\}$, from $D_{\text{mem}}$.
3.3.1 Hybrid Retrieval Strategy
To balance semantic understanding with keyword relevance, we employ a hybrid retrieval mechanism. We compute a relevance score for each trajectory $\tau\in D_{\text{mem}}$ as a weighted combination of dense and sparse retrieval scores:
$\text{Score}(\tau,I)=\alpha\cdot S_{\text{dense}}(\tau,I)+(1-\alpha)\cdot S_{\text{sparse}}(\tau,I),$ (3)
where $S_{\text{dense}}$ is the cosine similarity between the embeddings of $I$ and the final intent of $\tau$ (computed via FAISS), and $S_{\text{sparse}}$ is the BM25 lexical-match score. We retrieve the top-$K$ trajectories based on this score. Our experiments show $K=2$ yields an optimal trade-off (see Sec 4.3.3).
3.3.2 Memory Formatting
Raw trajectories are not directly consumable. Each retrieved trajectory $\tau_{k}\in M_{t}$ is processed into a structured, human-readable format. This involves summarizing each step into a tuple of {interface description, agent intent, action}. This transformation converts a raw log into an intuitive, step-by-step guide, making it an effective in-context learning example.
3.4 Stage III: Memory-Augmented Inference
In the final stage, the retrieved and formatted memories are used to guide a GUI agent, $\pi_{\text{agent}}$, which can be any off-the-shelf VLM. This highlights the plug-and-play nature of our framework.
At each inference step $t$, the agent’s behavior is conditioned on a structured prompt $P_{t}$, which integrates multiple sources of information:
$P_{t}=f(I,M_{t},H_{t},s_{t},E_{\text{sum}}(s_{t})),$ (4)
where $f$ is a template function, $I$ is the task instruction, $M_{t}$ are the formatted memories, $H_{t}$ is the current action history, $s_{t}$ is the screenshot, and $E_{\text{sum}}$ is a summarization model that provides a textual description of $s_{t}$. The agent then produces the next action $a_{t}\sim\pi_{\text{agent}}(P_{t})$. By providing explicit, successful examples ($M_{t}$), we effectively guide the agent’s reasoning process. The agent itself determines task completion by outputting a special ‘finish‘ action from its augmented action space $\mathcal{A}\cup\{\text{finish}\}$.
4 Experiments
In this section, we present a rigorous empirical evaluation of the EchoTrail-GUI framework. Our experiments are designed to answer three primary research questions: (1) Does EchoTrail-GUI outperform state-of-the-art GUI agents on established benchmarks? (2) What is the quantifiable contribution of each component within our framework? (3) How effective and efficient is our proposed self-exploration mechanism?
4.1 Experimental Settings
4.1.1 Environments and Baselines
We conduct our evaluation on two widely-adopted, interactive Android benchmarks: AndroidWorld [16], which features a diverse set of real-world applications, and AndroidLab [30], providing a more controlled environment for systematic testing.
Our framework’s performance is compared against a comprehensive suite of recent and state-of-the-art agents. This includes powerful proprietary models like GPT-4o and Gemini-1.5-Pro, as well as specialized open-source agents such as UGround [6] and UI-TARS [15]. To demonstrate the model-agnostic nature of EchoTrail-GUI, we implement our framework on two distinct backbones: a leading open-source model, Qwen2.5-VL-72B-Instruct, and a leading proprietary model, GPT-4o.
4.1.2 Evaluation Metrics
To ensure a comprehensive assessment, we adopt a multi-faceted set of metrics from the AndroidLab benchmark. SR, Sub-SR, RRR, and ROR stand for Success Rate, Sub-Goal Success Rate, Reversed Redundancy Ratio, and Reasonable Operation Ratio, respectively.
  • SR%:Success Rate. The primary metric, representing the percentage of tasks completed successfully.
  • Sub-SR%:Sub-Goal Success Rate. The percentage of correctly completed sub-goals, which measures fine-grained progress within a task.
  • RRR%:Reversed Redundancy Ratio. Measures the agent’s step efficiency. A higher RRR indicates that the agent took fewer redundant steps compared to an expert trajectory.
  • ROR%:Reasonable Operation Ratio. Measures the quality of individual actions by calculating the proportion of operations that are deemed “reasonable” or non-erroneous, indicating operational robustness.
For the AndroidWorld benchmark, we report the standard Success Rate.
4.1.3 Implementation Details
Our framework’s components are implemented with the following specifications:
  • Exploration Agent: Powered by Gemini 2.5 Flash. Maximum trajectory length is 30 steps.
  • Reward Model: Implemented with Gemini 2.5 Flash Lite, assigning a 1-5 score where the quality threshold $\theta_{\text{good}}$ is 4.
  • Inference Agent: We use Qwen2.5-VL-72B-Instruct and GPT-4o as backbones.
  • Summarization & Embedding: We utilize Qwen3-30B-Instruct-2507 for summaries and Qwen3-Embedding-4B for embeddings.
  • Retrieval: A hybrid FAISS and BM25 system retrieves $K=2$ memories, a choice justified by our sensitivity analysis in Section 4.3.3.
4.2 Main Results
We first evaluate the end-to-end performance of EchoTrail-GUI against all baselines. As a training-free, plug-and-play framework, our approach provides a substantial performance uplift to existing models by simply augmenting them with experiential memory. The detailed results are presented in Table 1 and Table 2.
4.2.1 Performance on AndroidWorld
As shown in Table 1, EchoTrail-GUI establishes itself as a leading framework on the AndroidWorld benchmark. When applied to a powerful proprietary model like GPT-4o, our framework yields substantial performance gains, significantly outperforming the base model by a wide margin. Notably, the resulting performance also surpasses other sophisticated augmentation methods, including those that require task-specific training.
This strong trend holds in the open-source category. EchoTrail-GUI markedly enhances the performance of Qwen2.5-VL-72B-Instruct. This enhancement allows the base model to match the performance of top fine-tuned open-source agents, yet it is achieved without any model retraining, validating the power and efficiency of our memory-driven approach.
4.2.2 Performance on AndroidLab
The results on AndroidLab, presented in Table 2, further corroborate these findings across a more comprehensive set of metrics. On the Qwen2.5-VL backbone, EchoTrail-GUI more than doubles the Success Rate and delivers consistent, significant improvements across all other evaluation dimensions, including sub-goal completion and operational efficiency. This indicates the agent becomes not only more successful but also more direct in its actions and robust in its operation.
A similar pattern is observed when augmenting GPT-4o, where our framework again provides a considerable boost to the Success Rate and all related metrics. By equipping powerful base models with a mechanism to learn from experience, EchoTrail-GUI consistently unlocks a new level of performance across diverse tasks and environments.
Table 1: Performance comparison on AndroidWorld benchmark. The table categorizes agents into Closed and Open Source, and indicates whether they operate without specific fine-tuning on GUI data (i.e., are ’Training-Free’). The best performance values are highlighted in bold. Our method is highlighted in gray.
[…]
Closed Source AppAgent [37] GPT-4o 14.9
Gemini [20] Gemini-1.5-Pro 22.8
Claude [2] Claude Computer-Use 27.9
UGround [6] GPT-4o + UGround 32.8
GPT-4o [1] GPT-4o 34.5
Ponder&Press [24] GPT-4o + Ponder&Press 34.5
Aguvis [32] GPT-4o + Aguvis 37.1
M3A [16] GPT-4o 40.5
ScaleTrack [8] GPT-4o + ScaleTrack 44.0
Aria-UI [34] GPT-4o + Aria-UI 44.8
URST [4] GPT-4o + URST Reflexion 46.6
AndroidGen [10] GPT-4o 46.8
GUI-explorer [28] GPT-4o 47.4
EchoTrail-GUI GPT-4o 51.7
Open Source Aguvis [32] Aguvis-72B 26.1
Qwen2.5-VL [3] Qwen2.5-VL-72B-Instruct 35.0
RAG-GUI [29] RAG-GUI-72B-RSF 45.7
UI-TARS [15] UI-TARS-72B-SFT 46.6
EchoTrail-GUI Qwen2.5-VL-72B-Instruct 46.6
Table 2: Performance comparison on AndroidLab benchmark. The table categorizes agents into Closed and Open Source, and indicates whether they operate without specific fine-tuning on GUI data (i.e., are ’Training-Free’). The best performance values are highlighted in bold. Our method is highlighted in gray.
[…]
Closed Source Gemini[19] Gemini-1.0 12.6 72.5 76.7 10.9
Claude [2] Claude-3-Opus 15.1 81.4 83.9 13.0
Gemini[20] Gemini-1.5-Pro 18.5 106.0 91.5 16.7
Claude [2] Claude-3.5-Sonnet 32.7 113.4 81.2 29.0
GPT-4o [1] GPT-4o 35.0 87.3 85.4 31.2
AutoGLM [12] AutoGLM 36.2
EchoTrail-GUI GPT-4o 50.7 97.9 88.5 48.1
Open Source UI-TARS [15] UI-TARS-72B 12.0 40.8 70.0 9.6
UI-TARS [15] UI-TARS-72B-ft 28.4 81.4 81.6 22.1
Qwen2.5-VL [3] Qwen2.5-VL-72B-Instruct 26.1 68.7 81.4 23.9
Qwen2-VL[22] Qwen2-VL-72B-Instruct-ft 29.3 84.5 90.2 24.6
Qwen2.5-VL [3] Qwen2.5-VL-72B-Instruct-ft 30.9 81.3 79.3 25.0
EchoTrail-GUI Qwen2.5-VL-72B-Instruct 41.1 89.4 92.1 37.5
4.3 Ablation and In-Depth Analysis
Having established the overall effectiveness of EchoTrail-GUI, we now dissect the framework to understand the contribution of its individual components and the dynamics of the self-exploration process.
4.3.1 Ablation Studies
To assess the contribution of each design choice, we perform a comprehensive ablation study on AndroidWorld using the Qwen2.5-VL-72B-Instruct backbone. As summarized in Table 3, we progressively remove components from the full EchoTrail-GUI framework and examine their impact on overall performance.
The full system provides a substantial improvement over the backbone model, demonstrating the effectiveness of integrating high-quality memory into the agent’s decision process. Among all factors, the removal of Critic-based Filtering leads to the most pronounced degradation. In this case, performance drops not only below that of the complete framework, but even below the backbone model operating without any memory at all, confirming a central premise of our approach: low-quality or noisy trajectories can mislead the agent and are therefore more harmful than omitting memory entirely. This underscores the necessity of a reliable critic to curate a trustworthy and semantically coherent memory base.
Eliminating the Hybrid Retrieval mechanism also results in a clear decline, indicating that retrieving contextually aligned trajectories is essential for providing meaningful guidance. Random or poorly matched examples fail to support reasoning and may introduce irrelevant or conflicting signals. Similarly, disabling Real-time Guidance during exploration reduces the quality and diversity of the collected experiences, leading to a weaker memory base and consequently less effective downstream inference.
Taken together, these ablations reveal that each component of EchoTrail-GUI plays a distinct yet complementary role: critic filtering ensures memory reliability, hybrid retrieval ensures relevance, and real-time guidance ensures memory quality and coverage. Their combined effect is necessary to realize the framework’s full potential.
Table 3: Ablation study on AndroidWorld. We start with the backbone model and show the performance of our full framework, followed by variants with specific components removed.
Method AndroidWorld
Easy Medium Hard Avg. SR (%)
Qwen2.5-VL-72B-Instruct 46.7 23.6 13.2 34.1
  w/o Critic-based Filtering 47.5 13.9 10.5 31.0
  w/o Hybrid Retrieval 60.7 20.8 13.2 40.5
  w/o Real-time Guidance 62.3 25.0 13.2 42.7
EchoTrail-GUI 65.6 30.6 15.8 46.6
4.3.2 Analysis of the Self-Exploration Mechanism
A core novelty of our work is the autonomous construction of the memory base. We analyze two key aspects of this process.
Quality and Diversity of Generated Trajectories.
A key claim of our work is that critic-guided self-exploration can autonomously generate a high-quality and diverse memory base. To validate this, we analyze the semantic properties of the trajectories in our generated EchoTrail-4K dataset. We use Sentence-BERT [17] to encode the final task intents from our explored trajectories and compare them against the ground-truth task instructions from the AndroidLab benchmark. The resulting embeddings are then visualized in two dimensions using UMAP [13], as shown in Figure 2.
The visualization reveals two crucial findings. First, the dense overlap between the generated tasks (blue) and the real tasks (orange) demonstrates a high degree of semantic alignment, confirming the realism of our exploration process. The agent is not generating random or nonsensical sequences, but rather tasks that are structurally and intentionally similar to human-defined goals. Second, the generated task space extends significantly beyond the boundaries of the original test set, filling in gaps and exploring new semantic regions. This confirms the diversity and effectiveness of our exploration strategy, showcasing its ability to build a comprehensive experience base that enhances the agent’s generalization capabilities.
Effectiveness of the Real-time Guidance Mechanism.
To validate the core hypothesis that our agent can learn from its own ongoing experiences, we analyze the effectiveness of the real-time guidance mechanism. We track the proportion of high-quality trajectories, as determined by our critic, generated over the course of exploration. This process is discretized into four sequential stages for several representative apps from the AndroidWorld environment. The results are illustrated in Figure 3.
The visualization reveals a clear and consistent upward trend across all tested applications, confirming that the agent’s exploration quality progressively improves over time. This provides strong evidence for the efficacy of our dual-database learning system. For complex applications like “OsmAnd” and “VLC”, the quality rate climbs by nearly 20 percentage points, showing rapid adaptation. For more intuitive applications like “Settings”, the framework helps consolidate an already high performance. This self-correcting behavior ensures the exploration process becomes increasingly efficient, populating the final memory database with reliable experiences.
Figure 2: UMAP visualization of task instructions from the AndroidLab test set (orange) and our self-explored trajectories (blue), demonstrating high semantic alignment and diversity.
Figure 3: The rate of high-quality trajectories generated across four sequential exploration stages on representative apps from AndroidWorld.
4.3.3 Analysis of Memory Injection
We further analyze the effect of the number of retrieved memories, $K$, on downstream performance. Using Qwen2.5-VL-72B-Instruct on AndroidWorld, we vary $K$ and measure the agent’s ability to exploit contextual guidance.
As shown in Figure 4, injecting a small set of semantically relevant trajectories consistently improves results relative to the memory-free baseline ($K=0$), confirming the value of targeted experiential context. Performance attains a maximum at a moderate retrieval size; beyond this point, additional trajectories yield diminishing returns and eventually degrade performance, a behavior we attribute to increased prompt length, conflicting advice, and contextual dilution.
These observations indicate an inherent trade-off in memory-augmented inference: targeted recall enhances decision-making, while excessive context can overwhelm the model and introduce noise. Based on this analysis, we select $K=2$ for all main experiments as a practical compromise between informativeness and contextual efficiency.
Figure 4: Sensitivity analysis of the number of injected memories ($K$) on AndroidWorld success rate.
5 Conclusion
In this work, we introduced EchoTrail-GUI, a unified framework that addresses the “digital amnesia” of contemporary GUI agents by enabling them to autonomously construct and leverage actionable memory. Through the integration of critic-guided self-exploration, lightweight yet semantically rich trajectory abstraction, and dynamic memory injection, the framework establishes an effective learn–remember–apply cycle that transforms stateless VLM-based agents into adaptive, experience-aware systems. Experiments on AndroidWorld and AndroidLab demonstrate substantial gains in success rate, sub-goal completion, and operational efficiency, while ablations highlight the critical role of high-quality trajectory filtering, hybrid retrieval, and real-time exploration guidance. Beyond performance improvements, EchoTrail-GUI shows that reliable operational memory can be built without manual supervision and that abstracted representations—comprising interface descriptions, inferred intentions, and executed actions—provide a scalable and transferable basis for contextual reasoning. Taken together, our findings position EchoTrail-GUI as a practical path toward GUI agents that continually improve through accumulated experience and offer a solid foundation for future work on lifelong interaction and multimodal memory system.